Shift summary - 12/1/99 Day shift

Julie Whitmore (jaws@fnal.gov)
Wed, 1 Dec 1999 15:49:45 -0600 (CST)

Shift Summary for day shift, 01/Dec/99

On Shift: Herman, Julie

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.5
Hours of E799 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Injection kicker problems and quench.

========================================================================
Runs ended this shift:
Run 15101 - 22.9M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)
UP*211-213,UPS071

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:
Relatively smooth E799 running. Spill structure a bit ratty because of QXR
problems. Accelerator down to work on injection kicker problem and then to
recover from a quench. After recovery, spill has 2Hz structure.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15101 7:00 Day shift takes over long E799 data run.

316 7:16 Ktevana error - missing FFFF word. KFREND.

8:00 Herman notices 3 spill periodicity to
bad spill structure. Phones MCR. They
believe QXR is responding to an unknown
instability in the machine. Experts are
looking into it and will keep us informed.

8:35 Onsplit request: UPH038 mounted on ktev2.

10:00 The usual: SDAQ RR71 -6V LV button is red.
ktsad.txt not updating. Probably a swap
space problem.

11:30 Onsplit request: UPF054 mounted on ktev1.

541-543 12:20 Zipped files in da_ktev runlog area. Also
cleaned up pipeline area.
MCR informs us that beam will disappear
for ~15min to work on F17 injection
kickers.

602 13:38 Beam disappears. End run w/ tapes ~80%
full. Mount new tapes.

15102 13:49 Start run 15102.

14:24 Abort due to injection quench. Some
permits were not reset properly after
kicker work.

36 14:35 Beam returns.

43 14:45 Herman calls MCR to complain about the 2Hz
spill structure. They are working on it.

14:54 "Lost frig permit in TeV. Est 10min."

Large Bry Air is off. Check unit outside.
Blower and pilot are off. Burnt smell
is detected. Ron and Leonard
paged. Ron wants us to keep an eye on
rib temps. Experts on route.

15:19 Bry Air reset not responding. Bld Mng is
calling duty electrician to open Bry Air
panel. Burnt odor not normal. Backup
system in KTeV hall operating normally.

63 15:25 Beam returns.

15:30 Wah is here. Where is his shift partner???

15:45 Page RickT. He is leaving meeting and will
be here in 7min. Damn that flagship
program!

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Bry Air unit has a broken relay. FESS is working on it.

The 2Hz structure is back. MCR is aware of it and is working on it. Might
want to make a follow-up call during the evening shift.